OpenStudy (surryyy):

The area of a square is 50 square meters. How long is each side? Round your answer to the nearest meter. Answer 2,500 100 50 7

Directrix (directrix):

The area of a square is the product of its length and its width or just the square of one of its sides because all sides of a square are congruent.

OpenStudy (surryyy):

so my answer would be 2,500?

OpenStudy (anonymous):

50^2=2500

OpenStudy (anonymous):

medal

OpenStudy (surryyy):

so was the final answer is 2500?

Directrix (directrix):

Wrong: 50^2=2500 --------------------- So, if x is a side of the square and the area is 50 square meters, then the equation you need to solve is this: x^2 = 50 So take the square root of both sides and post what you get.

OpenStudy (anonymous):

oh yeah so sqrt50=5 sqrt2

OpenStudy (anonymous):

only seven can be an answer

OpenStudy (anonymous):

7*7=49 49~~50

OpenStudy (surryyy):

x = 7

OpenStudy (anonymous):

yes

OpenStudy (surryyy):

Thnx